OpenStudy (anonymous):

A pair of equations is shown below: y = 2x − 1 y = 4x − 5 Part A: In your own words, explain how you can solve the pair of equations graphically. Write the slope and y-intercept for each equation that you will plot on the graph to solve the equations. (6 points) Part B: What is the solution to the pair of equations? (4 points)

OpenStudy (anonymous):

@Michele_Laino @thekefentse

OpenStudy (ddcamp):

You want a point (x,y) that is on both of those lines. This point will be where the two lines intersect. What would you do to find the intersection by graphing them?

OpenStudy (michele_laino):

hint: we can solve your system using the elimination method, namely, we make equal each other, the right sides of both equations

OpenStudy (michele_laino):

so we can write: \[2x - 1 = 4x - 5\]

OpenStudy (michele_laino):

please solve that equation for x

OpenStudy (anonymous):

x=2

OpenStudy (michele_laino):

that's right! nevertheless you need a graphical solution, I'm sorry for my misunderstand

OpenStudy (anonymous):

its ok. do you know how it can be put in as a graph?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

yea!

OpenStudy (michele_laino):

for example, let's consider the first equation: y=2x-1

OpenStudy (michele_laino):

if we set x=0, I get: y=2*0-1=...?

OpenStudy (anonymous):

-1

OpenStudy (michele_laino):

ok! so the firs line passes at point (0,-1)

OpenStudy (michele_laino):

furthermore, if I set y=0, then I can write: 0=2*x-1 what is x?

OpenStudy (anonymous):

1/2

OpenStudy (michele_laino):

ok! so our first line passes at points: A=(0,-1) and B=(1/2,0)

OpenStudy (anonymous):

yea thank you. medal of course. also i have one more question and ill stop bugging you.

OpenStudy (michele_laino):

now we can draw it as below: 

OpenStudy (anonymous):

I just saw this. Never got any notification :/

OpenStudy (anonymous):

hmm i put it there. i think. sorry.

OpenStudy (michele_laino):

now, let's consider the second equation: y=4x-5 again if I set x=0, I get: y=4*0-5=...?

OpenStudy (anonymous):

I see it. I'm new to this site, how/where do notifications work

OpenStudy (anonymous):

there will be a little sign that has a number over it and it will say someone tagged you.

OpenStudy (anonymous):

topish left corner of page

OpenStudy (anonymous):

ah! I see it go off now.

OpenStudy (anonymous):

y=-5

OpenStudy (michele_laino):

ok! so the second line passes at point P=(0,-5) now, I set y=0, so I can write this: 0=4x-5 what is x?

OpenStudy (anonymous):

5/4

OpenStudy (michele_laino):

ok! so the second line passes at points: P=(0,-5), and Q=(5/4,0)

OpenStudy (michele_laino):

here is the drawing of both lines: